Natural magnesium occurs as a mixture of three isotopes: Mg 24 12 at 78,8%, Mg 25 12 at 10,1% and Mg 26 12 at 11,2%. Calculate the atomic mass of a natural sample of magnesium.


1
Expert's answer
2021-12-23T10:49:48-0500

"A=0.788\cdot 24+0.101\cdot 25+0.112\cdot 26=24 .3."